OpenStudy (anonymous):

Please help: How do you simplify (2-2i)(-7-4i)

OpenStudy (anonymous):

How did you get 22?

OpenStudy (mrcoolguy):

-22+6i

OpenStudy (anonymous):

@mrcoolguy

OpenStudy (zenmo):

(2-2i)(-7-4i), use the FOIL method. [(2) x (-7)] + [(2) x (-4i)] + [(-2i) x (-7)] + [(-2i) x (-4i)] (-14) + (-8i) + (14i) + (8i^2) i^2=(-1), so (8i^2)= -8 -8i - 14 - 8 = -8i - 22